OpenStudy (agentnao):

y = 2 tan^3x I need to find an equation of the tangent line. I know you find the slope (derivative) and then what? How do I get the coordinates of the points?

OpenStudy (agentnao):

ANYONE WITH KNOWLEDGE OF DERIVATIVES: Your help would be greatly appreciated...

OpenStudy (anonymous):

The equation about what point?

OpenStudy (agentnao):

at x =3

OpenStudy (anonymous):

Okay so take the derivative as of that as you said. And substitute 3 into that equation.

OpenStudy (anonymous):

because the derivative represents the slope which is what you want for the tangent line.

OpenStudy (anonymous):

@AgentNao

OpenStudy (agentnao):

Oh! Yes, sorry! Okay, so I got the slope. And then? Is 3 the x coordinate and I just have to plug 3 into the original equation for the y value?

OpenStudy (anonymous):

Yep!

OpenStudy (agentnao):

OH! Crap. I'm sorry. Hang on. I just blended the information for two different problems here. The equation I gave you was right, but it's x = pi/4. I got the slope to be: 6 tan^2(x)sec^2(x), but I don't know how to solve it when plugging in pi/4 for x! How do you solve tan^2 of something?

OpenStudy (anonymous):

What's tan pi/4 ?

OpenStudy (agentnao):

1

OpenStudy (anonymous):

What's that squared? :P .

OpenStudy (agentnao):

But it's not tan (pi/4)^2, it's tan^2(pi/4)! They're not the same.

OpenStudy (anonymous):

That's the same thing as tan(pi/4) * Tan(pi/4) = 1*1 =1

OpenStudy (agentnao):

So I would normally use the product rule for that, but since u and v are the same in this case, it's just one! Okay. That makes sense. Thank you!

OpenStudy (anonymous):